What is the potential energi of an object where the mass is 4kg and the height is 8.5 meters


1
Expert's answer
2021-01-12T12:11:48-0500

By the definition of the potential energy, we have:


"P=mgh=4\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot 8.5\ m=333.2\ J."

Answer:

"P=333.2\ J."